# RADIO Amplified: Improved Baselines for Agglomerative Vision Foundation Models

### v1

@@ Abstract
Agglomerative models have recently emerged as a powerful approach to training vision foundation models, leveraging multi-teacher distillation from existing models such as CLIP, DINO, and SAM. This strategy enables the efficient creation of robust models, combining the strengths of individual teachers while significantly reducing computational and resource demands. In this paper, we thoroughly analyze state-of-the-art agglomerative models, identifying critical challenges including resolution mode shifts, teacher imbalance, idiosyncratic teacher artifacts, and an excessive number of output tokens. To address these issues, we propose several novel solutions: multi-resolution training, mosaic augmentation, and improved balancing of teacher loss functions. Specifically, in the context of Vision Language Models, we introduce a token compression technique to maintain high-resolution information within a fixed token count. We release our top-performing models, available in multiple scales (-B, -L, -H, and -g), alongside inference code and pretrained weights.
@@ 1 Introduction
Figure 1: Overview of the main contributions: From left to right, we introduce a multi-resolution training framework that enables our student model to maintain accuracy across all resolutions; we propose using token merging to better retain fine-grained details while merging uninformative tokens; we improve upon the AM-RADIO baseline across all benchmarks, with significant gains on VLM tasks.
The rise of specialized vision foundation models has created a need for methods to consolidate knowledge from multiple models into a single one. SAM-CLIP [39] addresses this challenge by combining SAM [19] and CLIP [27] to integrate capabilities from both. In another approach, AM-RADIO [30] introduces label-free knowledge distillation from multiple teacher models, enabling knowledge transfer without direct supervision. UNIC [32] adds intermediate teacher matching projectors and dynamic teacher selection. Theia [33] aims to facilitate robot learning by distilling insights from multiple vision teachers. Meanwhile, PHI-S [29] studies the importance of normalizing the distinct teacher distributions to simplify their balancing. Relatedly, Eagle [34] leverages a mixture of vision encoders to achieve inference-time knowledge aggregation within the context of vision-language models (VLMs).
@@
  • Generating multi-resolution features for diverse applications: Vision models support various applications requiring different feature resolutions, from image captioning to dense segmentation. A VFM that can flexibly produce features at any resolution could serve multiple tasks, reducing the need for separate models, and opening up new opportunities.
In our analysis of existing agglomerative models, we study and propose a fix for the notable “mode switch” phenomenon in AM-RADIO, where feature distributions shift significantly based on input resolution (Section 3.1). Specifically, low-resolution inputs yield DINO-like features, while high-resolution inputs produce SAM-like features. We trace this behavior to the student learning different teachers at different resolutions during training. In section 4.2, we introduce a solution to stabilize these mode switches, achieving strong resolution robustness and improved quality scaling.
Armed now with a vision encoder that works best at high resolution, we next look toward integrating it into VLMs. In the case of downstream applications with vision-language models, a common pitfall is the number of output tokens/features. By processing images in high resolution, most methods will return more image tokens and will result in quadratic complexity of attention in VLMs. High resolution processing is important, and in order to preserve this information, we propose to apply ToMeSD [3] in section 4.7 to decouple the vision encoder resolution from the number of patches used by the VLM.
Our main contributions are as follows:
@@
Model Goal ImageNet1K Segmentation Probe3D[11] Vision-Language (VILA[20]) SAM [19]
@@
AM-RADIO-L* Baseline 78.35 84.01 50.03 82.63 59.94 57.01 15.6 24.19 28.88 25.60 73.77
+ multi-res Eliminate modes 81.21 84.09 52.84 82.39 61.04 59.44 16.6 33.99 29.02 29.40 75.49
- OpenAICLIP + SigLIP Better VLM 81.01 84.68 52.95 84.7 60.1 64.42 25.28 45.80 30.60 36.00 74.19
+ ViT-H Bigger backbone 82.51 85.81 53.97 85.7 62.5 65.88 25.96 49.74 35.17 40.90 76.14
+ Token Merging Improve VLM - - - - - 69.74 30.40 52.33 36.24 42.90 -
Table 1: Ablation Results. For Probe3D, Depth and Surface Normals metrics are averaged over the buckets defined in the paper. For SAM COCO we use the “instance all” bucket. Each incremental change we introduce leads to improved metrics. *Our reproduction.
@@ 2.1 Baseline Model
Following AM-RADIO[30], our model consists of a Vision Transformer (ViT)[10] backbone, including CPE for multi-resolution support. We use DFN CLIP [13], OpenAI CLIP [27], DINOv2-g-reg [9] and SAM-H [19] as teachers. For each teacher to distill from, we augment the backbone with an adaptor for the summarization token, and an adaptor for the patch tokens. Our adaptor is a 2-layer MLP with a LayerNorm[1] and a GeLU[16] in between. We train the student using teacher features generated by inferring images of the DataComp1B[14] dataset. We train the model with a batch size of 1024+128 for 600k iterations. Training is split into two concurrent partitions: one partition trains the student at a resolution of $432^{2}$ against CLIP and DINOv2 teachers with batch size 1024, while the other partition trains the student at a resolution of $1024^{2}$ against SAM with a batch size of 128.
@@ 2.2 Evaluation Framework
  • Pixel-level semantic segmentation: We linearly probe the frozen backbone features for semantic segmentation.
  • Instance segmentation: We evaluate our model on the COCO [18] dataset.
@@ 3.1 Achieving Multi-Resolution Robustness
Figure 2: Zero-Shot Accuracy and teacher matching fidelity metrics as a function of input resolution. AM-RADIO, and our “Multi-Stage Stage 3” model exhibit mode switching behavior. “Multi-Stage Stage 2” doesn’t exhibit the behavior, owing to all teachers being trained solely at low resolution, however quality at high resolution quickly degrades. “Multi-Resolution” training fixes the mode switch, and allows the model to be strong across a larger range of resolutions compared to Stage 2.
@@
Figure 3: Visualizations of patch token features for different input resolutions. We use PCA to project patch tokens into a 3D-space representing RGB colors. Top: From left to right: input image, visualizations of DINOv2, AM-RADIO at $256^{2}$, $384^{2}$, $512^{2}$, $768^{2}$ and $1024^{2}$ resolutions, and SAM. Bottom: RADIO-AMP (ours) features. As resolution increases, AM-RADIO features look increasingly similar to those of SAM, even exhibiting VitDet windowing artifacts. RADIO-AMP features are much more self-consistent and less noisy. More samples can be found in the Appendix.
@@ 4.1.2 Tiling
Figure 4: Qualitative comparison of tiling and high-resolution inference. Tiling emulates high-res SigLIP[44] inference. Top: From left to right: input, SigLIP features using 1, $2\times 2$ and $4\times 4$ and $5\times 5$ tiles. Bottom: Single-pass inference using RADIO-AMP (ours) at equivalent image sizes. SigLIP features vary significantly across different tiling arrangements.
Figure 5: Visualizations of SigLIP and RADIO-AMP model features before and after ToMe compression. From left to right: input image; PCA visualization of features; merged tokens; PCA visualization of unmerged tokens. RADIO-AMP is better able to compress background regions, leaving more tokens available for semantically rich features.
@@ 4.2 Multi-Resolution Training
RADIO-AMP-B 0.102 0.168
RADIO-AMP-L 0.102 0.165
RADIO-AMP-H 0.119 0.193
RADIO-AMP-L (tiling) N/A 0.369
@@ 4.3 Mosaic Augmentation
To improve efficiency when training a student at a resolution $\leq 512^{2}$ against SAM, we can instead create a mosaic of $k\times k$ images, with $k=\left\lfloor\frac{1024}{x}\right\rfloor$ and $x$ being the student resolution, resulting in a single $1024^{2}$ image. We then perform SAM inference on this mosaic and extract $k^{2}$ individual feature maps to train the student. Mosaic augmentation includes padding as needed to maximize efficiency. For example, to train a student at $432^{2}$ resolution, we can create a $2\times 2$ mosaic with 80-pixel padding around each image. Figure 7 shows sample mosaic augmentations under 256 and 432 student resolutions. Qualitatively, we observe cleaner features after applying mosaic augmentation, which we believe is due to the increased diversity in image positions, helping to reduce positional encoding artifacts. We also show SAM’s PCA features for these mosaic images in figures A5 and A4 in the appendix.
@@ 4.4 Teacher Loss Balancing
PHI-S [29] highlights the significant variations in activation magnitudes among vision foundation models, observing that SAM’s activations tend to overshadow those of CLIP and DINOv2 models. We adopt the PCA-Hadamard Isotropic Standardization (PHI-S) method, achieving improved balance among teacher losses. PHI-S rotates teacher activations to evenly distribute variance across all channels and then scales them to obtain unit variance. This process can be easily reversed by projecting the student activations back into each teacher’s original feature space. PHI-S enhances training stability and overall benchmark performance. For a given teacher feature map $\mathbf{X}$ with embedding size $C$, PHI-S applies the following transformation:
@@ 4.7 Token Compression
Inspired by ToMeSD[3], we evaluate the use of bipartite soft matching to merge similar tokens. We apply strided partitioning to ensure that each image region retains some representation in the compressed features. For evaluation, we track merged token indices, enabling us to unmerge tokens and measure the reconstruction error and inform hyperparameter selection. The method is illustrated in Figure 8. While ToMe recommends using attention keys for matching, we found that using the features themselves improves reconstruction error. Table A9 of the appendix summarizes our findings.
@@
We further note that token merging can be applied to other vision encoders and demonstrate its application on SigLIP. Interestingly, token merging yields a greater relative improvement on RADIO-AMP than on SigLIP, a finding that aligns with the feature visualizations in Figure 5 and the measured reconstruction error in Table A9. We hypothesize that SigLIP’s higher feature noise makes clustering more error prone.
@@ 4.8 Feature Selection
Table 7: Study on the effect of RADIO-AMP-H feature selection. For “dense” aggregation, the numbers in brackets indicate the range of layers from which the average is calculated. Top: Semantic segmentation, depth estimation, surface normals estimation. Bottom: VILA benchmarks. Pixel-level tasks exhibit a clear preference for more layers and non-linear heads, while VLM tasks seem mostly neutral to this choice.
4.9 SigLIP Teacher
Building on previous related work ([38], [20]), we replace OpenAI-CLIP [27] with SigLIP [44] and observe significant improvements in VLM tasks, as shown in Table 1.

### v2

# RADIOv2.5: Improved Baselines for Agglomerative Vision Foundation Models

@@ Abstract
Agglomerative models have recently emerged as a powerful approach to training vision foundation models, leveraging multi-teacher distillation from existing models such as CLIP, DINO, and SAM. This strategy enables the efficient creation of robust models, combining the strengths of individual teachers while significantly reducing computational and resource demands. In this paper, we thoroughly analyze state-of-the-art agglomerative models, identifying critical challenges including resolution mode shifts, teacher imbalance, idiosyncratic teacher artifacts, and an excessive number of output tokens. To address these issues, we propose several novel solutions: multi-resolution training, mosaic augmentation, and improved balancing of teacher loss functions. Specifically, in the context of Vision Language Models, we introduce a token compression technique to maintain high-resolution information within a fixed token count. We release our top-performing variants at multiple scales (-B, -L, -H, and -g), along with inference code and pretrained weights.
@@ 1 Introduction
Figure 1: Overview of the main contributions: From left to right, we introduce a multi-resolution training framework that enables our student model to maintain accuracy across all resolutions; we propose using token merging to better retain fine-grained details while merging uninformative tokens; our RADIOv2.5 models improve upon the RADIOv2.1 baseline across all benchmarks, with significant gains on VLM tasks.
The rise of specialized Vision Foundation Models (VFMs) has created a need for methods to consolidate knowledge from multiple models into a unified model. SAM-CLIP [45] addresses this challenge by combining SAM [23] and CLIP [33] to integrate capabilities from both. In another approach, AM-RADIO [36] introduces label-free knowledge distillation from multiple teacher models, enabling knowledge transfer without direct supervision. UNIC [38] adds intermediate teacher matching projectors and dynamic teacher selection. Theia [39] aims to facilitate robot learning by distilling insights from multiple vision teachers. Meanwhile, PHI-S [35] studies the importance of normalizing the distinct teacher distributions to simplify their balancing. Relatedly, Eagle [40] leverages a mixture of vision encoders to achieve inference-time knowledge aggregation within the context of Vision-Language Models (VLMs).
@@
  • Generating multi-resolution features for diverse applications: Vision models support various applications requiring different feature resolutions, from image captioning to dense segmentation. A VFM that flexibly produces features at any resolution could serve multiple tasks, reducing the need for separate models and unlocking new opportunities.
In our analysis of existing agglomerative models, we study and propose a fix for the notable “mode switch” phenomenon in AM-RADIO, where feature distributions shift significantly based on input resolution (Section 3.1). Specifically, low-resolution inputs yield DINO-like features, while high-resolution inputs produce SAM-like features. We trace this behavior to the student learning from different teachers at different resolutions during training. In section 4.2, we introduce a solution to stabilize these mode switches, achieving strong resolution robustness and improved quality scaling.
Armed now with a vision encoder that works best at high resolution, we next look toward integrating it into VLMs. In the case of downstream applications with vision-language models, a common pitfall is the number of output tokens/features. By processing images in high resolution, most methods will return more image tokens and will result in quadratic complexity of attention in VLMs. High resolution processing is important, and in order to preserve this information, we propose applying ToMeSD [4] in section 4.8 to decouple the vision encoder resolution from the number of patches used by the VLM.
Our main contributions, illustrated in Figure 1, are:
@@
Configuration Goal ImageNet1K Segmentation Probe3D[13] Vision-Language (VILA[24]) SAM [23]
@@
$\mathcal{A}$: RADIOv2.1-L* Baseline 78.35 84.01 50.03 82.63 59.94 57.01 15.6 24.19 28.88 25.60 73.77
$\mathcal{B}$: $\mathcal{A}$ + multi-res Eliminate modes 81.21 84.09 52.84 82.39 61.04 59.44 16.6 33.99 29.02 29.40 75.49
$\mathcal{C}$: $\mathcal{B}$ OpenAICLIP + SigLIP Better VLM 81.01 84.68 52.95 84.7 60.1 64.42 25.28 45.80 30.60 36.00 74.19
$\mathcal{D}$: $\mathcal{C}$ + ViT-H Bigger backbone 82.51 85.81 53.97 85.7 62.5 65.88 25.96 49.74 35.17 40.90 76.14
$\mathcal{E}$: $\mathcal{D}$ + Token Merging Improve VLM - - - - - 69.74 30.40 52.33 36.24 42.90 -
Table 1: Ablation Results. For Probe3D, Depth and Surface Normals metrics are averaged over the buckets defined in the paper. For SAM COCO we use the “instance all” bucket. Each incremental change we introduce leads to improved metrics. *We use a ViT-L instead of the ViT-H used in the original AM-RADIO [36] paper.
@@ 2 Background
2.1 Knowledge Agglomeration
The assumption underlying knowledge agglomeration is that multiple foundation models exist, each capable of extracting diverse and meaningful representations from a wide range of internet-sourced images. Furthermore, it assumes that knowledge from these models can be distilled into a single agglomerative model.
Let $x$ be the input data. The student’s shared backbone produces a feature representation:
$\mathbf{z}=f(x)=\Big{[}z_{s},\,z_{p}^{(1)},\,z_{p}^{(2)},\,\dots,\,z_{p}^{(N)}\Big{]},$ (1)
where:
  • $z_{s}\in\mathbb{R}^{d}$ is the summary token,
  • $z_{p}^{(i)}\in\mathbb{R}^{d}$ for $i=1,\dots,N$ are the patch tokens,
  • $d$ is the student’s embedding dimension.
For each teacher $t$, the student’s model includes two adaptor heads:
  1. 1.
    $g_{s}^{(t)}:\mathbb{R}^{d}\to\mathbb{R}^{d_{t}}$, which projects the summary token $z_{s}$ into the teacher’s embedding space.
  2. 2.
    $g_{p}^{(t)}:\mathbb{R}^{d}\to\mathbb{R}^{d_{t}}$, which projects each patch token $z_{p}^{(i)}$.
These adaptor heads are typically implemented as simple multi-layer perceptrons (MLPs) that can adjust the feature dimension as required. For example, one may define:
$g^{(t)}(z)=\mathrm{MLP}^{(t)}(z)=\sigma\Big{(}W^{(t)}z+b^{(t)}\Big{)},$ (2)
where $\sigma(\cdot)$ is a non-linear activation function, and $W^{(t)}$ and $b^{(t)}$ are learnable parameters.
Thus, for each teacher $t$, the projected features are:
Summary Projection: $\hat{z}_{s}^{(t)}=g_{s}^{(t)}(z_{s}),$ (3)
Patch Projection: $\hat{z}_{p}^{(t)}=g_{p}^{(t)}(z_{p}).$ (4)
These projected features are then used in a knowledge distillation process that encourages the student to mimic the representations of each teacher, effectively aggregating diverse knowledge into a single agglomerative model.
The training objective is to align the student features with the corresponding teacher features. To achieve this, we define a loss function that computes an aggregate measure of similarity between the teacher and student features. Let $\mathcal{L}_{t}$ denote the loss for teacher $t$. This loss can be defined as:
$\mathcal{L}_{t}=\ell_{s}\Big{(}\hat{z}_{s}^{(t)},z_{s}^{(t)}\Big{)}+\sum_{i=1}^{N}\ell_{p}\Big{(}\hat{z}_{p}^{(t,i)},z_{p}^{(t,i)}\Big{)},$ (5)
where:
  • $\ell_{\cdot}(\cdot,\cdot)$ is a similarity or distance metric (e.g., mean squared error or cosine similarity loss)
  • $\ell_{s}$, the summary loss objective, need not be the same as $\ell_{p}$, the patch loss objective
  • $z_{s}^{(t)}$ and $z_{p}^{(t,i)}$ are the summary and patch features extracted from teacher $t$.
The overall loss function aggregates the losses from all teachers:
$\mathcal{L}=\sum_{t}\lambda_{t}\,\mathcal{L}_{t},$ (6)
where $\lambda_{t}$ are weighting factors that balance the contribution of each teacher’s loss.
@@ 2.2 Baseline Model
Following AM-RADIO[36], our baseline model consists of a Vision Transformer (ViT)[12] backbone, including CPE [22] for multi-resolution support. We use DFN CLIP [16], OpenAI CLIP [33], DINOv2-g-reg [10] and SAM-H [23] as teachers. For each teacher to distill from, we augment the backbone with an adaptor for the summarization token, and an adaptor for the patch tokens. Our adaptor is a 2-layer MLP with a LayerNorm[1] and a GeLU[19] in between. We train the student using teacher features generated by inferring images of the DataComp1B[17] dataset. For faster iterations, we replicate AM-RADIO using a ViT-L backbone instead of the original ViT-H backbone and refer to it as RADIOv2.1-L from this point forward. We train the model with a batch size of 1024+128 for 600k iterations. Training is split into two concurrent partitions: one partition trains the student at a resolution of $432^{2}$ against CLIP and DINOv2 teachers with batch size 1024, while the other partition trains the student at a resolution of $1024^{2}$ against SAM with a batch size of 128. This baseline is referred to as $\mathcal{A}$ in Table 1.
@@ 2.3 Evaluation Framework
  • Pixel-level semantic segmentation: We linearly probe the frozen backbone features for semantic segmentation on both ADE20k [53] and Pascal VOC [14].
  • Instance segmentation: We evaluate our model on the COCO [22] dataset using the protocol from EfficientViT [26], and our learned SAM adaptor.
@@
  • Additional dense multi-task probing on Pascal Context [31] and NYUDv2 [32] following the setup in MLoRE [50].
@@ 3.1 Achieving Multi-Resolution Robustness
Figure 2: Zero-Shot Accuracy and teacher matching fidelity metrics as a function of input resolution. AM-RADIO (aka RADIOv2.1, config $\mathcal{A}$), and our “Multi-Stage Stage 3” model exhibit mode switching behavior. “Multi-Stage Stage 2” doesn’t exhibit the behavior, owing to all teachers being trained solely at low resolution, however quality at high resolution quickly degrades. “Multi-Resolution” (config $\mathcal{B}$) training fixes the mode switch, and allows the model to be strong across a larger range of resolutions compared to Stage 2.
@@
Figure 3: Visualizations of patch token features across input resolutions. We use PCA to map patch tokens to RGB colors. Top: Input image, followed by visualizations of DINOv2, RADIOv2.1 (config $\mathcal{A}$) at $256^{2}$, $512^{2}$, $1024^{2}$ resolutions, and SAM. Bottom: RADIOv2.5-H (config $\mathcal{D}$). As resolution increases, RADIOv2.1 features resemble SAM’s, even showing VitDet windowing artifacts, while RADIOv2.5 remains more self-consistent and less noisy.
@@ 4.1.2 Tiling
Figure 4: Qualitative comparison of tiling and high-resolution inference. Tiling emulates high-res SigLIP inference. Top: From left to right: input, SigLIP features using 1, $2\times 2$ and $4\times 4$ tiles. Bottom: Single-pass inference using RADIOv2.5 (ours) at equivalent image sizes. SigLIP features vary significantly across different tiling arrangements.
@@ 4.2 Multi-Resolution Training
Multi-resolution training is illustrated in Figure 5, which depicts four training regimes based on whether the teachers and/or students process low- or high-resolution images. When both the student and teachers use either low- or high-resolution images, no particular challenge arises. However, training a high-resolution student against a low-resolution teacher requires downscaling the student’s features to match those of the teacher. Conversely, training a low-resolution student against a high-resolution teacher necessitates a careful augmentation technique, as described in Section 4.3.
@@
RADIOv2.5-B 0.102 0.168
RADIOv2.5-L 0.102 0.165
RADIOv2.5-H 0.119 0.193
RADIOv2.5-L (tiling) N/A 0.369
@@ 4.3 Mosaic Augmentation
To improve efficiency when training a student at a resolution $\leq 512^{2}$ against SAM, we can instead create a mosaic of $k\times k$ images, with $k=\left\lfloor\frac{1024}{x}\right\rfloor$ and $x$ being the student resolution, resulting in a single $1024^{2}$ image. We then perform SAM inference on this mosaic and extract $k^{2}$ individual feature maps to train the student. Mosaic augmentation includes padding as needed to maximize efficiency. For example, to train a student at $432^{2}$ resolution, we can create a $2\times 2$ mosaic with 80-pixel padding around each image. Figure 6 shows sample mosaic augmentations under 256 and 432 student resolutions. Qualitatively, we observe cleaner features after applying mosaic augmentation, which we believe is due to the increased diversity in image positions, helping to reduce positional encoding artifacts. We also show SAM’s PCA features for these mosaic images in figures A4 and A5.
@@ 4.4 Teacher Loss Balancing
PHI-S [35] highlights the significant variations in activation magnitudes among vision foundation models, observing that SAM’s activations tend to overshadow those of CLIP and DINOv2 models. We adopt the PCA-Hadamard Isotropic Standardization (PHI-S) method, achieving improved balance among teacher losses. PHI-S rotates teacher activations to evenly distribute variance across all channels and then scales them to obtain unit variance. This process can be easily reversed by projecting the student activations back into each teacher’s original feature space, or even by modifying the final linear projection of the adaptor MLP after training. PHI-S enhances training stability and overall benchmark performance. For a given teacher feature map $\mathbf{X}$ with embedding size $C$, PHI-S applies the following transformation:
@@ 4.6 Partitioning
4.7 SigLIP Teacher
Building on previous work ([44], [24]), we replace OpenAI-CLIP [33] with SigLIP [52], defining this as our configuration $\mathcal{C}$. Our choice is validated by the significant improvements observed in VLM tasks, as shown in Table 1.
@@ 4.8 Token Compression
Figure 7: Visualizations of SigLIP and RADIOv2.5-H model features before and after ToMe compression. From left to right: input image; PCA visualization of features; merged tokens; PCA visualization of unmerged tokens. RADIOv2.5 is better able to compress background regions, leaving more tokens available for semantically rich features.
@@
Inspired by ToMeSD[4], we evaluate the use of bipartite soft matching to merge similar tokens. We apply strided partitioning to ensure that each image region retains some representation in the compressed features. For evaluation, we track merged token indices, enabling us to unmerge tokens and measure the reconstruction error and inform hyperparameter selection. The method is illustrated in Figure 8.
While ToMe recommends using attention keys for matching, we found that using the features directly reduces reconstruction error. Table A10 in the appendix summarizes our findings. Another slight departure from ToMe’s recommendation is that we apply merging at the output of the vision encoder rather than incrementally. As shown in Table A4, this approach yields an average improvement of +1.4 on VLM benchmarks.
@@
We further note that token merging can be applied to other vision encoders and demonstrate its application on SigLIP. Interestingly, token merging yields a greater relative improvement on RADIOv2.5 than on SigLIP, a finding that aligns with the feature visualizations in Figure 7 and the measured reconstruction error in Table A10. We hypothesize that SigLIP’s higher feature noise makes clustering more error prone.
@@ 4.9 Feature Selection
Table 7: Study on the effect of RADIOv2.5-H feature selection. For “dense” aggregation, the numbers in brackets indicate the range of layers from which the average is calculated. Top: Semantic segmentation, depth estimation, surface normals estimation. Bottom: VILA benchmarks. Pixel-level tasks exhibit a clear preference for more layers and non-linear heads, while VLM tasks seem mostly neutral to this choice.